Buoyancy aids in watersports are tested by adding increasing weights until they sink. A


club has two sets of buoyancy aids. One set is two years old, and should support weights


according to ๐‘(6.0,0.64)kg, the other set is five years old and should support weights of


๐‘(4.5, 1.0) kg. all the aids are tested and any which are unable to spport at least 5kg are


thrown out.


a. If there are 24 two-year aids, how many are still usable?


b. If there are 32 five-year old aids how many are still usable?

1
Expert's answer
2022-05-16T23:32:16-0400

a.


"P(X>5)=1-P(Z\le\dfrac{5-6}{\sqrt{0.64}})"

"=1-P(Z\le-1.25)=0.89435"

"0.89435(24)=21"

21 buoyancy aids.


b.


"P(X>5)=1-P(Z\le\dfrac{5-4.5}{\sqrt{1.0}})"

"=1-P(Z\le0.5)=0.30854"

"0.30854(32)=10"

10 buoyancy aids.